OpenStudy (anonymous):

0.3x - 0.2y = 4 and 0.4 = 0.3 y= -1

OpenStudy (anonymous):

What are we doing with these?

OpenStudy (anonymous):

There might be one too many equal signs there. Did you copy the problem correctly?

OpenStudy (anonymous):

Alright - so, to get started, first identify which variable would be easier to cancel by way of elimination. Unfortunately, with these, there is no variable that is easier to cancel than any other, so you can just pick one. I'm going to go with "y". So let's look at causing the y's to cancel. 0.4x + 0.3y = -1 0.3x - 0.2y = 4

OpenStudy (anonymous):

ok i tried doing that but was more confused

OpenStudy (anonymous):

Now, the full name for this way of solving the system is "Elimination by Addition" So what we're trying to do is make it so the "y" values have the same coefficient (number in front of them) by we want them to have opposite signs.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So first, to make things easy, let's get rid of those decimals by multiplying both equations by 10. 10(0.4x + 0.3y = -1) 10(0.3x - 0.2y = 4)

OpenStudy (anonymous):

That would leave us with: 4x + 3y = -10 3x - 2y = 40

OpenStudy (anonymous):

From here, let's multiply the top equation by "-2" and the bottom equation by "3". This will cause the y variables to cancel when we add these two equations together.

OpenStudy (anonymous):

4(-2)x + 3(-2)y = (-2)-10 3(3)x - 2(3)y = (3)40

OpenStudy (anonymous):

-8x + -6y = 20 9x - 6y = 120

OpenStudy (anonymous):

Now, we can add them together.

OpenStudy (anonymous):

That would leave us with 1x=140

OpenStudy (anonymous):

got that so far

OpenStudy (anonymous):

So we found that x=140 Now, you can plug that value into either equation for the "x" and solve for "y"

OpenStudy (anonymous):

so then y would be -230 over -6

OpenStudy (anonymous):

taking 1st equation 0.4x + 0.3y = -1 0.4x = -0.3y - 1 x = -(0.3y + 1) / 0.4 now put this value in 2nd equation ... 0.3x - 0.2y = 4 [-0.3(0.3y+1)] / 0.4 - 0.2y = 4 -0.09y/0.4 - 0.3/0.4 -0.2y = 4 from this extract value of y then put in 1st eqn , u will get of 'x'

OpenStudy (anonymous):

4(140)+3y=-10 560+3y=-10 3y=-570 y=-190

OpenStudy (anonymous):

So the solution would be (140, -190)

OpenStudy (anonymous):

ok thank you this help me out

OpenStudy (anonymous):

Happy to help! :)

OpenStudy (accessdenied):

It's always a good gesture to give the helper a "Good Answer" medal to show your appreciation! :)

OpenStudy (anonymous):

solution is (-4.15, 2.2)

OpenStudy (anonymous):

I'm sorry about that. I found the mistake: From here, let's multiply the top equation by "2" and the bottom equation by "3". This will cause the y variables to cancel when we add these two equations together. 4(2)x + 3(2)y = (2)-10 3(3)x - 2(3)y = (3)40 8x+6y=-20 9x-6y=120 So when we add them, we get: 17x=100 x=100/17 Then, to get "y" plug (100/17) in for "x" in either equation. We'll do it for both of them just to make sure we're right this time. 8(100/17)+6y=-20 800/17+6y=-20 6y=-1140/17 y=-190/17 If we plug our x value of 100/17 into the other equation we get: 9(100/17)-6y=120 900/17-6y=120 -6y=1140/17 y=-190/17 So we know we got the right answer this time. ;) So the solution should be : (100/17, -190/17) PS: You were right, that one was tricky!!